OpenStudy (anonymous):

a square piece of cardboard is to be formed into a box to transport pizzas. The box is formed by cutting 2-inch squares corners from the cardboard and folding them up. If the volume of the box is 512 in^3, what are the dimensions of the cardboard? Is this a volume or area problem?

OpenStudy (amistre64):

\[V_{box}=height*width*length\]

OpenStudy (amistre64):

since its a square; the width and length are the same

OpenStudy (amistre64):

the problem is an area problem

OpenStudy (amistre64):

512 = 2(side-2)^2 512/2 = (side-2)^2 sqrt(512/2) = side-2 2 + sqrt(512/2) = side

OpenStudy (amistre64):

that comes to each side was originally 18 inches before being cut out

OpenStudy (amistre64):

but i gotta wonder just how an open box transports pizzas

OpenStudy (anonymous):

so add the 2-inch for the corners that are folded up make the dimensions 20 x 20

OpenStudy (anonymous):

very good question :)

OpenStudy (amistre64):

no; the equation factors that 2 inches in to get 18 inches for a starting value

OpenStudy (anonymous):

so 18 x 18 dimensions

OpenStudy (amistre64):

to dbl chk; 18-2 = 16 16^2 = 256 256*2 = something close to 512 :)

OpenStudy (amistre64):

that dont compute right does it...

OpenStudy (amistre64):

it should be 4 inches from each side; 2 from each corner

OpenStudy (anonymous):

yes bc it's a square so 2 x 2=4

OpenStudy (amistre64):

512 = 2(side-4)^2 512/2 = (side-4)^2 sqrt(512/2) = side-4 4 + sqrt(512/2) = side

OpenStudy (amistre64):

whaddayaknow; that makes it 20 per side lol

OpenStudy (anonymous):

lol....ty

OpenStudy (amistre64):

yw :)